The small magnets each of magnetic moment 10 Am^2 are placed end-on position 0.1m apart from their centres.The force actjng between them is?

Grade:12th Pass

1 Answers

Arun
4 years ago
Dear student
 
 
Formula for force BT parallel placed magnet f=Mo/4π(6m1m2/r4)
No=4π×10^-7 therefore f=10^-7(6×10×10/(0.1^4)=0.6N